# CoInfra: A Large-Scale Cooperative Infrastructure Perception System and Dataset in Adverse Weather

### v1

@@ III-D Delay-Aware Synchronization Protocol
  • Adaptive Fusion Window: For each synchronized time anchor, the system waits for incoming data from all nodes within an adaptively determined fusion window (typically $3\sigma$ of the estimated latency). If a message from a node is delayed beyond this window or dropped, its data is excluded from the global fusion at that cycle.
  • Delay-Aware Data Fusion and Correction: All available sensor data within the fusion window is combined using appropriate fusion strategies (early, deep, or late fusion). For late-arriving data (after the window), our framework supports confidence adjustment or direct post-fusion using delay-adaptive motion models to update the global perception map.

### v2

@@ III-D Delay-Aware Synchronization Protocol
  • Adaptive Fusion Window: For each synchronized time anchor, the system waits for incoming data from all nodes within an adaptively determined fusion window (typically $3\sigma$ or $4\sigma$ of the estimated latency). If a message from a node is delayed beyond this window or dropped, its data is excluded from the global fusion at that cycle.
  • Delay-Aware Data Fusion and Correction: All available sensor data within the fusion window is combined using appropriate fusion strategies (early, deep, or late fusion). For late-arriving data (after the window), our framework supports confidence adjustment or direct post-fusion using delay-adaptive motion models to update the global perception results.
@@
III-E Sensor Node and Cloud Level Perception
The sensor node and cloud level perception adopted in CoInfra builds on the previous work [25].
The local perception pipeline in each sensor node follows a modular design and consists of the following stages:
  • 2D Object Detection: Images from the RGB cameras are processed using deep neural networks, such as custom-trained YOLO variants[26], to identify objects as 2D bounding boxes in the camera frame.
  • Region of Interest (ROI) Filtering: Point clouds are filtered spatially to remove irrelevant background points using the HD map and global calibration information.
  • Hierarchical Clustering with Scanning Pattern Awareness: Remaining points are first clustered within individual scanning lines, then grouped across lines using a custom distance metric. It accelerates computation and improves clustering robustness, especially in weather-affected scenarios. This results in a set of 3D clusters representing potential objects.
  • Multi-modal Fusion with Ground Contact Features: Detected 2D bounding boxes are associated with 3D point cloud clusters by leveraging LiDAR-camera projection and ground-contacting cues. This enables robust semantic association of 3D clusters and then generates accurate 3D bounding boxes using cluster and semantic knowledge.
  • Object Tracking: Sequential detections are temporally linked using a multi-object tracking algorithm to maintain object identities across frames and smooth detections.
The above pipeline is implemented on the edge (e.g., Jetson Orin NX), providing low-latency inference and reducing bandwidth by transmitting only structured outputs (detections and tracks). Depending on node resources, the perception stack can be further extended with more advanced early or deep fusion methods.
At the cloud level, the CoInfra system performs delay-aware global fusion and tracking by aggregating structured outputs from all sensor nodes. The main steps include:
  • Time-Synchronized Data Fusion: The cloud collects node-level outputs that are temporally aligned using global time anchors. Incoming data is classified as either normal (arriving within an expected window) or late (arriving after the window due to abnormal delays).
  • Confidence-Aware Integration: Normal data is fused across nodes using weighted averaging for robust object association. Late-arriving data is assigned reduced confidence and either incorporated with caution or used to update existing tracks post hoc.
  • Delay-Adaptive Multi-Object Tracking: The system maintains global tracks for all detected objects, continuously updating trajectories by combining new detections with existing tracks.
  • Motion Model-Based Delay Correction: The system applies motion prediction models to estimate object positions at the current fusion time, compensating for delay and ensuring accurate, up-to-date scene understanding.